# A search for Galactic post-AGB stars in Gaia DR3††thanks: Tables A6 and A7 are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A

### v1

@@ Acknowledgements.
This work has made use of data from the European Space Agency (ESA) Gaia mission, processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. This research has made use of the Simbad database and the Aladin sky atlas, operated at CDS, Strasbourg, France. The authors have also made use of the VOSA software, developed under the Spanish Virtual Observatory project supported by the Spanish MINECO through grant PID2020-112949GB-I00, and partially funded by the European Union’s Seventh Framework Programme (FP7-SPACE-2013-1) for research, technological development and demonstration under grant agreement no. 60674. Funding from Spanish Ministry project PID2021-122842OB-C22, Xunta de Galicia ED431B 2021/36 and PDC2021-121059-C22 is acknowledged by the authors. We also acknowledge support from CIGUS-CITIC, funded by Xunta de Galicia and the European Union (FEDER Galicia 2014-2020 Program) through grant ED431G 2019/01. E.V. acknowledges support from the ’DISCOBOLO’ funded by the Spanish Ministerio de Ciencia, Innovación y Universidades under grant PID2021-127289NB-I00. M.M. and E.V. acknowledge support from the cooperation agreement between the IAC and the Fundación Jesús Serra for visiting grants. AM acknowledges support from the ACIISI, Gobierno de Canarias and the European Regional Development Fund (ERDF) under grant with reference PROID2020010051 as well as from the State Research Agency (AEI) of the Spanish Ministry of Science and Innovation (MICINN) under grant PID2020-115758GB-I00.References
@@ Appendix A Data Tables
Num | Simbad Name | Gaia DR3 ID | RA | Dec | G mag | Spectral Type | Reference | Flag | |
@@
2 | 1307 | 1199 | 1436 | 0.50 | 6500 | T3(1) | 4.11 | 4.02 | 4.06 | 3.14 | disc | - | |
7 | 1909 | 1870 | 1947 | 3.85 | 4250 | T3(2) | 4.34 | 4.32 | 2.39 | 2.79 | 1.22 | stellar | H |
14 | 6639 | 5922 | 7516 | 1.75 | 16000 | T3(3) | 3.75 | 3.63 | 3.74 | 1.33 | 3.62 | stellar | - |
15 | 2271 | 2160 | 2400 | 4.75 | 3750 | T1 | 3.61 | 3.56 | 2.32 | 1.12 | stellar | H | |
18 | 3467 | 3159 | 3934 | 0.20 | 20000 | T3(4) | 3.55 | 3.44 | 3.47 | 0.93 | 18.14 | shell | H |
19 | 2839 | 2711 | 2962 | 1.60 | 6500 | T3(5) | 3.62 | 3.58 | 3.41 | 1.12 | 5.61 | shell | - |
21 | 4734 | 4150 | 5414 | 0.30 | 21000 | T3(6) | 3.61 | 3.47 | 3.95 | 0.95 | 16.54 | shell | H |
24 | 9965 | 9131 | 11304 | 0.90 | 22000 | T3(7) | 4.30 | 4.19 | 4.15 | 3.81 | 1.10 | stellar | (*) |
25 | 5862 | 5008 | 7121 | 0.73 | 23000 | T3(8) | 4.40 | 4.21 | 4.02 | 3.00 | 9.15 | stellar | H |
26 | 4982 | 4299 | 6049 | 1.00 | 18000 | T3(9) | 4.24 | 4.06 | 3.88 | 3.63 | 0.99 | stellar | - |
27 | 4718 | 4214 | 5732 | 3.30 | 6000 | T2 | 4.12 | 3.95 | 3.98 | 3.16 | disc | - | |
29 | 3858 | 3644 | 4143 | 2.15 | 20000 | T3(6) | 4.43 | 4.37 | 3.09 | 4.00 | 0.69 | shell | - |
30 | 3421 | 3226 | 3635 | 3.14 | 9250 | T2 | 3.87 | 3.81 | 3.66 | 1.98 | 3.36 | shell | - |
31 | 4340 | 4100 | 4599 | 1.97 | 9000 | T2 | 3.77 | 3.72 | 3.63 | 1.42 | 3.94 | shell | - |
33 | 2822 | 2677 | 2982 | 1.78 | 20000 | T3(6) | 4.02 | 3.97 | 3.45 | 2.84 | 2.14 | shell | - |
34 | 5021 | 4607 | 5683 | 1.00 | 8500 | T2 | 3.76 | 3.65 | 3.72 | 1.38 | 3.82 | shell | - |
35 | 10125 | 9128 | 11379 | 2.25 | 6750 | T2 | 3.67 | 3.56 | 3.27 | 0.97 | 9.06 | stellar | H |
37 | 10312 | 8988 | 12281 | 11750 | T2 | 3.45 | 3.28 | 3.70 | 13.83 | shell | H | ||
39 | 2916 | 2712 | 3196 | 0.80 | 7250 | T3(10) | 3.92 | 3.84 | 4.04 | 2.27 | 1.49 | stellar | (*) |
40 | 2099 | 1991 | 2209 | 2.00 | 7250 | T3(11) | 3.94 | 3.89 | 3.37 | 2.40 | 1.63 | shell | H |
41 | 1104 | 1075 | 1139 | 1.00 | 4250 | T3(12) | 4.16 | 4.13 | 4.08 | 3.22 | disc | - | |
42 | 1788 | 1722 | 1845 | 0.82 | 7250 | T3(13) | 3.95 | 3.92 | 3.96 | 2.42 | 1.66 | disc | - |
43 | 1567 | 1503 | 1632 | 0.25 | 8250 | T3(14) | 3.54 | 3.50 | 3.46 | 0.93 | 9.14 | stellar | H |
45 | 5130 | 4530 | 5923 | 0.70 | 7000 | T3(13) | 4.40 | 4.27 | 4.40 | 3.00 | 9.14 | disc | H |
47 | 1708 | 1647 | 1782 | 1.00 | 8250 | T3(11) | 3.74 | 3.71 | 3.63 | 1.32 | 3.57 | shell | - |
49 | 3101 | 2895 | 3388 | 0.60 | 6750 | T3(11) | 4.51 | 4.44 | 4.62 | 3.00 | 9.14 | stellar | H |
51 | 1921 | 1830 | 2016 | 0.75 | 6000 | T3(11) | 3.88 | 3.83 | 3.71 | 2.07 | 0.28 | shell | H |
53 | 1807 | 1726 | 1891 | 0.40 | 9750 | T3(15) | 3.38 | 3.34 | 3.18 | 9.67 | stellar | - | |
54 | 6999 | 5822 | 8254 | 1.40 | 7000 | T3(16) | 4.56 | 4.37 | 4.51 | 3.00 | 9.14 | disc | H |
58 | 2310 | 2165 | 2481 | 1.80 | 8000 | T3(11) | 3.89 | 3.83 | 3.58 | 2.12 | 1.34 | shell | - |
59 | 1390 | 1344 | 1439 | 0.35 | 10000 | T3(17) | 3.57 | 3.54 | 3.41 | 1.07 | 5.54 | stellar | H (*) |
60 | 1410 | 1356 | 1465 | 2.80 | 5250 | T3(11) | 3.94 | 3.90 | 3.44 | 2.35 | 0.58 | shell | - |
62 | 5416 | 5040 | 5792 | 2.19 | 14000 | T3(18) | 3.51 | 3.44 | 3.50 | 10.68 | shell | - | |
63 | 3685 | 3461 | 3942 | 1.80 | 22000 | T3(19) | 4.15 | 4.09 | 2.88 | 3.32 | 0.79 | shell | - |
64 | 4817 | 4587 | 5042 | 3.00 | 5500 | T3(11) | 3.47 | 3.43 | 3.47 | 8.54 | shell | - | |
72 | 5022 | 4555 | 5658 | 3.35 | 9000 | T3(18) | 3.99 | 3.88 | 3.37 | 2.62 | 1.88 | shell | - |
75 | 9082 | 8261 | 10230 | 2.70 | 7500 | T3(11) | 3.88 | 3.78 | 3.61 | 2.08 | 1.29 | shell | - |
98 | 6605 | 5703 | 7346 | 2.25 | 6750 | T2 | 3.38 | 3.25 | 3.12 | 9.14 | disc | - | |
99 | 2082 | 2026 | 2148 | 3.75 | 14000 | T3(20) | 3.91 | 3.88 | 3.31 | 2.23 | 1.45 | shell | - |
102 | 6830 | 5887 | 8163 | 6.75 | 9750 | T3(18) | 3.69 | 3.51 | 3.10 | 1.20 | 8.33 | shell | - |
111 | 10863 | 9455 | 12375 | 1.40 | 15000 | T3(21) | 4.23 | 4.09 | 3.70 | 2.58 | 1.83 | stellar | H |
112 | 7934 | 6777 | 9233 | 1.75 | 19000 | T3(6) | 4.41 | 4.25 | 3.95 | 3.00 | 1.84 | shell | H |
113 | 6437 | 5646 | 7505 | 1.60 | 16000 | T3(22) | 3.44 | 3.29 | 3.20 | 11.32 | shell | H | |
114 | 5354 | 4801 | 6044 | 2.10 | 19000 | T3(23) | 4.01 | 3.89 | 3.40 | 2.77 | 2.07 | shell | - |
116 | 7318 | 6417 | 8628 | 1.05 | 22000 | T3(6) | 4.12 | 3.96 | 3.27 | 2.26 | 1.55 | shell | H |
117 | 5974 | 5146 | 7444 | 1.00 | 5750 | T3(24) | 3.22 | 3.00 | 3.38 | 0.90 | 7.96 | stellar | H |
121 | 7513 | 6290 | 9363 | 0.30 | 17000 | T3(25) | 3.34 | 3.11 | 3.48 | 19.81 | stellar | H | |
127 | 6147 | 5365 | 6967 | 1.20 | 6250 | T3(13) | 3.46 | 3.32 | 3.34 | 8.96 | disc | - | |
128 | 2317 | 2237 | 2410 | 3.50 | 5750 | T3(26) | 3.73 | 3.69 | 3.34 | 1.24 | 8.85 | disc | - |
129 | 4556 | 4246 | 4929 | 1.35 | 6750 | T3(27) | 3.34 | 3.26 | 3.12 | 9.14 | disc | - | |
132 | 10720 | 9413 | 12080 | 1.75 | 8500 | T2 | 4.48 | 4.35 | 4.07 | 3.00 | 9.15 | disc | H |
133 | 1684 | 1619 | 1769 | 1.00 | 6250 | T3(11) | 4.02 | 3.97 | 3.92 | 2.71 | 1.92 | disc | - |
135 | 1813 | 1707 | 1923 | 2.40 | 6000 | T3(28) | 4.49 | 4.43 | 3.63 | 4.00 | 4.05 | disc | - |
@@
137 | 7253 | 6109 | 8872 | 0.70 | 6000 | T2 | 3.40 | 3.17 | 3.46 | 8.29 | stellar | H | |
138 | 2369 | 2250 | 2507 | 2.25 | 6500 | T2 | 3.46 | 3.40 | 2.94 | 8.96 | stellar | - | |
139 | 4682 | 4285 | 5128 | 1.75 | 6250 | T2 | 3.86 | 3.77 | 3.54 | 1.18 | 7.64 | shell | H |
141 | 3948 | 3726 | 4264 | 2.00 | 6250 | T3(29) | 3.37 | 3.31 | 3.35 | 8.96 | disc | - | |
142 | 727 | 714 | 743 | 1.00 | 6750 | T3(11) | 3.47 | 3.45 | 3.25 | 9.14 | shell | - | |
143 | 4188 | 3841 | 4631 | 1.20 | 5000 | T3(27) | 3.67 | 3.58 | 3.65 | 1.18 | 7.13 | disc | - |
144 | 4269 | 3980 | 4580 | 2.00 | 7250 | T2 | 4.10 | 4.03 | 3.91 | 3.13 | 0.67 | disc | - |
145 | 4795 | 4442 | 5245 | 1.50 | 6250 | T2 | 3.41 | 3.32 | 3.11 | 8.96 | stellar | - | |
146 | 4842 | 4502 | 5201 | 2.25 | 5500 | T2 | 3.58 | 3.51 | 3.28 | 1.08 | 4.78 | stellar | - |
147 | 8617 | 7584 | 9718 | 1.70 | 4500 | T2 | 3.34 | 3.21 | 3.40 | 6.83 | stellar | - | |
148 | 5122 | 4745 | 5526 | 10.25 | 7000 | T3(30) | 4.13 | 4.06 | 3.48 | 3.22 | disc | - | |
152 | 2116 | 1995 | 2245 | 2.80 | 5500 | T3(31) | 3.40 | 3.34 | 3.27 | 8.54 | stellar | - | |
153 | 4404 | 4113 | 4693 | 2.00 | 7500 | T3(27) | 4.42 | 4.35 | 4.36 | 4.00 | 9.14 | disc | - |
156 | 1535 | 1468 | 1606 | 1.80 | 6750 | T2 | 3.64 | 3.60 | 3.38 | 1.14 | 6.95 | shell | - |
157 | 2348 | 2251 | 2450 | 2.80 | 4250 | T3(32) | 3.41 | 3.37 | 3.27 | 5.58 | stellar | - |

### v2

@@ Acknowledgements.
This work has made use of data from the European Space Agency (ESA) Gaia mission, processed by the Gaia Data Processing and Analysis Consortium (DPAC). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. This research has made use of the Simbad database and the Aladin sky atlas, operated at CDS, Strasbourg, France. The authors have also made use of the VOSA software, developed under the Spanish Virtual Observatory project supported by the Spanish MINECO through grant PID2020-112949GB-I00, and partially funded by the European Union’s Seventh Framework Programme (FP7-SPACE-2013-1) for research, technological development and demonstration under grant agreement no. 60674. This research was funded by the Spanish Ministry of Science MCIN / AEI / 10.13039 / 501100011033 and the European Union Next Generation programme EU/PRTR through the coordinated grant PID2021-122842OB-C22 and the Horizon Europe [HORIZON-CL4-2023-SPACE-01-71] SPACIOUS project, Grant Agreement no. 101135205. Additionally, it is co-financed by the EU through the FEDER Galicia 2021-27 operational programme, Ref. ED431G 2023/01. E.V. acknowledges support from the ’DISCOBOLO’ funded by the Spanish Ministerio de Ciencia, Innovación y Universidades under grant PID2021-127289NB-I00. M.M. and E.V. acknowledge support from the cooperation agreement between the IAC and the Fundación Jesús Serra for visiting grants. AM acknowledges support from the ACIISI, Gobierno de Canarias and the European Regional Development Fund (ERDF) under grant with reference PROID2020010051 as well as from the State Research Agency (AEI) of the Spanish Ministry of Science and Innovation (MICINN) under grant PID2020-115758GB-I00.References
@@ Appendix A Data Tables
Num | Simbad Name | Gaia EDR3 ID | RA | Dec | G mag | Spectral Type | Reference | Flag | |
@@
2 | 1307 | 1199 | 1436 | 0.50 | 6500 | T3(1) | 4.11 | 4.02 | 4.20 | 3.14 | disc | - | |
7 | 1909 | 1870 | 1947 | 3.85 | 4250 | T3(2) | 4.34 | 4.32 | 4.37 | 2.79 | 1.22 | stellar | H |
14 | 6639 | 5922 | 7516 | 1.75 | 16000 | T3(3) | 3.75 | 3.63 | 3.88 | 1.33 | 3.62 | stellar | - |
15 | 2271 | 2160 | 2400 | 4.75 | 3750 | T1 | 3.61 | 3.56 | 3.66 | 1.12 | stellar | H | |
18 | 3467 | 3159 | 3934 | 0.20 | 20000 | T3(4) | 3.55 | 3.44 | 3.67 | 0.93 | 18.14 | shell | H |
19 | 2839 | 2711 | 2962 | 1.60 | 6500 | T3(5) | 3.62 | 3.58 | 3.66 | 1.12 | 5.61 | shell | - |
21 | 4734 | 4150 | 5414 | 0.30 | 21000 | T3(6) | 3.61 | 3.47 | 3.75 | 0.95 | 16.54 | shell | H |
24 | 9965 | 9131 | 11304 | 0.90 | 22000 | T3(7) | 4.30 | 4.19 | 4.42 | 3.81 | 1.10 | stellar | (*) |
25 | 5862 | 5008 | 7121 | 0.73 | 23000 | T3(8) | 4.40 | 4.21 | 4.58 | 3.00 | 9.15 | stellar | H |
26 | 4982 | 4299 | 6049 | 1.00 | 18000 | T3(9) | 4.24 | 4.06 | 4.43 | 3.63 | 0.99 | stellar | - |
27 | 4718 | 4214 | 5732 | 3.30 | 6000 | T2 | 4.12 | 3.95 | 4.29 | 3.16 | disc | - | |
29 | 3858 | 3644 | 4143 | 2.15 | 20000 | T3(6) | 4.43 | 4.37 | 4.49 | 4.00 | 0.69 | shell | - |
30 | 3421 | 3226 | 3635 | 3.14 | 9250 | T2 | 3.87 | 3.81 | 3.93 | 1.98 | 3.36 | shell | - |
31 | 4340 | 4100 | 4599 | 1.97 | 9000 | T2 | 3.77 | 3.72 | 3.83 | 1.42 | 3.94 | shell | - |
33 | 2822 | 2677 | 2982 | 1.78 | 20000 | T3(6) | 4.02 | 3.97 | 4.08 | 2.84 | 2.14 | shell | - |
34 | 5021 | 4607 | 5683 | 1.00 | 8500 | T2 | 3.76 | 3.65 | 3.88 | 1.38 | 3.82 | shell | - |
35 | 10125 | 9128 | 11379 | 2.25 | 6750 | T2 | 3.67 | 3.56 | 3.78 | 0.97 | 9.06 | stellar | H |
37 | 10312 | 8988 | 12281 | 11750 | T2 | 3.45 | 3.28 | 3.63 | 13.83 | shell | H | ||
39 | 2916 | 2712 | 3196 | 0.80 | 7250 | T3(10) | 3.92 | 3.84 | 4.01 | 2.27 | 1.49 | stellar | (*) |
40 | 2099 | 1991 | 2209 | 2.00 | 7250 | T3(11) | 3.94 | 3.89 | 3.99 | 2.40 | 1.63 | shell | H |
41 | 1104 | 1075 | 1139 | 1.00 | 4250 | T3(12) | 4.16 | 4.13 | 4.19 | 3.22 | disc | - | |
42 | 1788 | 1722 | 1845 | 0.82 | 7250 | T3(13) | 3.95 | 3.92 | 3.98 | 2.42 | 1.66 | disc | - |
43 | 1567 | 1503 | 1632 | 0.25 | 8250 | T3(14) | 3.54 | 3.50 | 3.58 | 0.93 | 9.14 | stellar | H |
45 | 5130 | 4530 | 5923 | 0.70 | 7000 | T3(13) | 4.40 | 4.27 | 4.53 | 3.00 | 9.14 | disc | H |
47 | 1708 | 1647 | 1782 | 1.00 | 8250 | T3(11) | 3.74 | 3.71 | 3.77 | 1.32 | 3.57 | shell | - |
49 | 3101 | 2895 | 3388 | 0.60 | 6750 | T3(11) | 4.51 | 4.44 | 4.58 | 3.00 | 9.14 | stellar | H |
51 | 1921 | 1830 | 2016 | 0.75 | 6000 | T3(11) | 3.88 | 3.83 | 3.93 | 2.07 | 0.28 | shell | H |
53 | 1807 | 1726 | 1891 | 0.40 | 9750 | T3(15) | 3.38 | 3.34 | 3.42 | 9.67 | stellar | - | |
54 | 6999 | 5822 | 8254 | 1.40 | 7000 | T3(16) | 4.56 | 4.37 | 4.74 | 3.00 | 9.14 | disc | H |
58 | 2310 | 2165 | 2481 | 1.80 | 8000 | T3(11) | 3.89 | 3.83 | 3.95 | 2.12 | 1.34 | shell | - |
59 | 1390 | 1344 | 1439 | 0.35 | 10000 | T3(17) | 3.57 | 3.54 | 3.60 | 1.07 | 5.54 | stellar | H (*) |
60 | 1410 | 1356 | 1465 | 2.80 | 5250 | T3(11) | 3.94 | 3.90 | 3.98 | 2.35 | 0.58 | shell | - |
62 | 5416 | 5040 | 5792 | 2.19 | 14000 | T3(18) | 3.51 | 3.44 | 3.58 | 10.68 | shell | - | |
63 | 3685 | 3461 | 3942 | 1.80 | 22000 | T3(19) | 4.15 | 4.09 | 4.22 | 3.32 | 0.79 | shell | - |
64 | 4817 | 4587 | 5042 | 3.00 | 5500 | T3(11) | 3.47 | 3.43 | 3.51 | 8.54 | shell | - | |
72 | 5022 | 4555 | 5658 | 3.35 | 9000 | T3(18) | 3.99 | 3.88 | 4.09 | 2.62 | 1.88 | shell | - |
75 | 9082 | 8261 | 10230 | 2.70 | 7500 | T3(11) | 3.88 | 3.78 | 3.97 | 2.08 | 1.29 | shell | - |
98 | 6605 | 5703 | 7346 | 2.25 | 6750 | T2 | 3.38 | 3.25 | 3.51 | 9.14 | disc | - | |
99 | 2082 | 2026 | 2148 | 3.75 | 14000 | T3(20) | 3.91 | 3.88 | 3.94 | 2.23 | 1.45 | shell | - |
102 | 6830 | 5887 | 8163 | 6.75 | 9750 | T3(18) | 3.69 | 3.51 | 3.87 | 1.20 | 8.33 | shell | - |
111 | 10863 | 9455 | 12375 | 1.40 | 15000 | T3(21) | 4.23 | 4.09 | 4.38 | 2.58 | 1.83 | stellar | H |
112 | 7934 | 6777 | 9233 | 1.75 | 19000 | T3(6) | 4.41 | 4.25 | 4.57 | 3.00 | 1.84 | shell | H |
113 | 6437 | 5646 | 7505 | 1.60 | 16000 | T3(22) | 3.44 | 3.29 | 3.60 | 11.32 | shell | H | |
114 | 5354 | 4801 | 6044 | 2.10 | 19000 | T3(23) | 4.01 | 3.89 | 4.13 | 2.77 | 2.07 | shell | - |
116 | 7318 | 6417 | 8628 | 1.05 | 22000 | T3(6) | 4.12 | 3.96 | 4.29 | 2.26 | 1.55 | shell | H |
117 | 5974 | 5146 | 7444 | 1.00 | 5750 | T3(24) | 3.22 | 3.00 | 3.43 | 0.90 | 7.96 | stellar | H |
121 | 7513 | 6290 | 9363 | 0.30 | 17000 | T3(25) | 3.34 | 3.11 | 3.58 | 19.81 | stellar | H | |
127 | 6147 | 5365 | 6967 | 1.20 | 6250 | T3(13) | 3.46 | 3.32 | 3.61 | 8.96 | disc | - | |
128 | 2317 | 2237 | 2410 | 3.50 | 5750 | T3(26) | 3.73 | 3.69 | 3.78 | 1.24 | 8.85 | disc | - |
129 | 4556 | 4246 | 4929 | 1.35 | 6750 | T3(27) | 3.34 | 3.26 | 3.42 | 9.14 | disc | - | |
132 | 10720 | 9413 | 12080 | 1.75 | 8500 | T2 | 4.48 | 4.35 | 4.61 | 3.00 | 9.15 | disc | H |
133 | 1684 | 1619 | 1769 | 1.00 | 6250 | T3(11) | 4.02 | 3.97 | 4.06 | 2.71 | 1.92 | disc | - |
135 | 1813 | 1707 | 1923 | 2.40 | 6000 | T3(28) | 4.49 | 4.43 | 4.55 | 4.00 | 4.05 | disc | - |
@@
137 | 7253 | 6109 | 8872 | 0.70 | 6000 | T2 | 3.40 | 3.17 | 3.64 | 8.29 | stellar | H | |
138 | 2369 | 2250 | 2507 | 2.25 | 6500 | T2 | 3.46 | 3.40 | 3.53 | 8.96 | stellar | - | |
139 | 4682 | 4285 | 5128 | 1.75 | 6250 | T2 | 3.86 | 3.77 | 3.95 | 1.18 | 7.64 | shell | H |
141 | 3948 | 3726 | 4264 | 2.00 | 6250 | T3(29) | 3.37 | 3.31 | 3.43 | 8.96 | disc | - | |
142 | 727 | 714 | 743 | 1.00 | 6750 | T3(11) | 3.47 | 3.45 | 3.49 | 9.14 | shell | - | |
143 | 4188 | 3841 | 4631 | 1.20 | 5000 | T3(27) | 3.67 | 3.58 | 3.77 | 1.18 | 7.13 | disc | - |
144 | 4269 | 3980 | 4580 | 2.00 | 7250 | T2 | 4.10 | 4.03 | 4.16 | 3.13 | 0.67 | disc | - |
145 | 4795 | 4442 | 5245 | 1.50 | 6250 | T2 | 3.41 | 3.32 | 3.51 | 8.96 | stellar | - | |
146 | 4842 | 4502 | 5201 | 2.25 | 5500 | T2 | 3.58 | 3.51 | 3.65 | 1.08 | 4.78 | stellar | - |
147 | 8617 | 7584 | 9718 | 1.70 | 4500 | T2 | 3.34 | 3.21 | 3.47 | 6.83 | stellar | - | |
148 | 5122 | 4745 | 5526 | 10.25 | 7000 | T3(30) | 4.13 | 4.06 | 4.21 | 3.22 | disc | - | |
152 | 2116 | 1995 | 2245 | 2.80 | 5500 | T3(31) | 3.40 | 3.34 | 3.47 | 8.54 | stellar | - | |
153 | 4404 | 4113 | 4693 | 2.00 | 7500 | T3(27) | 4.42 | 4.35 | 4.48 | 4.00 | 9.14 | disc | - |
156 | 1535 | 1468 | 1606 | 1.80 | 6750 | T2 | 3.64 | 3.60 | 3.68 | 1.14 | 6.95 | shell | - |
157 | 2348 | 2251 | 2450 | 2.80 | 4250 | T3(32) | 3.41 | 3.37 | 3.46 | 5.58 | stellar | - |